OpenStudy (anonymous):

The second term in a geometric sequence is 20. The fourth term in the same sequence is 45/4, or 11.25. What is the common ratio in this sequence?

OpenStudy (ilovecake):

hold on.

OpenStudy (anonymous):

ok

OpenStudy (ilovecake):

an=a×rn−1 a is the 1st term and r is the common ratio. so in your question you know 2nd term 20=ar2−1...or.....20=ar and the 4th term 11.25=ar4−1....or....11.25=ar3 the 4th term can be written as ar times r^2 since r * r^2 = r^3 so you can substitute then 2nd term 11.25=ar×r2....or........11.25=20×r2 now you just need to solve for r.

OpenStudy (ilovecake):

.75

OpenStudy (anonymous):

thank you

OpenStudy (ilovecake):

np